One of the roots of the equation y= -2x^2+kx+(k+2) is located at x=3
A) determine the value of k
B) determine the other root of this equation

Answer by htmentor(1343)

One of the roots of the equation y= -2x^2+kx+(k+2) is located at x=3
A) determine the value of k
B) determine the other root of this equation
To determine k, we substitute x=3 into the equation -2x^2 + kx + (k+2) = 0
This gives -2(3)^2 + 3k + k + 2 = 0
Solving for k gives k=4
So the equation is y = -2x^2 + 4x + 6
To get the other root, we know that one of the factors of this quadratic is x-3.
From this we can determine the other factor:
(x-3)(-2x-2) = 0
This gives x=-1